OpenStudy (anonymous):

domain and range of the function in interval notation f(x)=sqrtx-1

OpenStudy (anonymous):

\[f(x)=\sqrt{x-1}\]

OpenStudy (anonymous):

you cannot have negative squareroot, so set \[x-1 \ge0\], then solve for x, so we'll have \[x \ge1 \]. so domain is anything greater than 1, [1,infinity)

OpenStudy (anonymous):

i think domain is \[(1,\infty) and the range is [-1,\infty]\]

OpenStudy (anonymous):

and the range is going to be ??

OpenStudy (anonymous):

0 to infinity. why u think range is from -1 to infinity?

OpenStudy (anonymous):

because of the problem being x-1 I thought that would make it -1 to infinity but as you see I am wrong and do not get this at all

OpenStudy (anonymous):

would the thing be [ or( before the 0

OpenStudy (anonymous):

that's have nothing to do with range, range is more about y. What's the lowest y values?

OpenStudy (anonymous):

[0, infinity), because 0 is included, you can take sqrt[0] right?

OpenStudy (anonymous):

oh I am going to be so glad when today is over, today is my last class for my aa and it is driving me crazy. I still have 21 questions to go

OpenStudy (anonymous):

lol, goodluck